# Analysis of Clustering and Degree Index in Random Graphs and Complex Networks

### v1

@@ 4 Degree index in Erdős-Rényi graphs
Let us conclude the section with the proof of Proposition 4.1 regarding independent binomial random variables. Although this result is standard, we include the details for the sake of completeness.
@@
Proof of Proposition 4.1 (i) Let $X_{n},Y_{n}$ be independent binomial random variables with parameters $n$ and $1/2$. We claim that $\mathbb{E}|X_{n}-Y_{n}|=\frac{n\binom{2n}{n}}{2^{2n}}$. For this purpose, letting $Z_{n}$ be an independent binomial random variable with parameters $n$ and $1/2$, observe that
@@
(ii) Let $Z_{1},Z_{2}$ be independent standard normal random variables. We denote the Wasserstein distance between distributions by $d_{W}$. Now, letting $\mu_{n}=\mathbb{E}[X_{n}]$ and $\sigma_{n}^{2}=\mathrm{Var}(X_{n})$, recall that
@@
5 Monte Carlo simulations for other random graph models
The calculations for degree and clustering indices become more involved when one leaves the framework of Erdős-Rényi random graphs and deal with other models. In this section we focus on three other random graph models, and do a Monte Carlo simulation study for these. The models we will be interested in are the Watts-Strogatz model, Barabási-Albert model and lastly, the random regular graphs. Degree index in random regular graphs is already zero, and one is only interested in the clustering index for this special case.
5.1 Random graph models and implementation
Here, instead of providing a detailed technical background on the random graph models of interest, we will briefly discuss the practical aspects of the implementation of the Monte Carlo simulations. In particular, we will focus on the selection of parameters so that the experiments are done with matching edge densities for distinct random graph models. Before moving further let us note that we used the NetworkX module in Python for all the simulations done below. Now let us go over the models under study, keeping in mind that we are willing to have a given edge density $p_{*}\in(0,1)$ in our simulations.
1. Erdős-Rényi model. The function erdos_renyi_graph which has parameters $n$ and $p$ as discussed in previous sections produces an Erdős-Rényi graph in NetworkX. One just chooses $p=p_{*}$ in order to satisfy the edge density criteria.
2. Watts-Strogatz model. The model is introduced by D. J. Watts and S. Strogatz in 1998, and it catches certain common properties of real life complex networks by having the small world property, high clustering and short average path lengths [23]. The description of watts_strogatz_graph function in NetworkX reference page is as follows: First create a ring over $n$ nodes. Then each node in the ring is joined to its $k$ nearest neighbors (or $k-1$ neighbors if is odd). Then shortcuts are created by replacing some edges as follows: for each edge $(u,v)$ in the underlying “$n$-ring with $k$ nearest neighbors” with probability $p$ replace it with a new edge $(u,w)$ with uniformly random choice of existing node $w$.
The number of edges in Watts-Strogatz model is constant and is given by $\frac{nk}{2}$. Thus in order to have the edge density $p_{*}$, one needs to have $p_{*}=\frac{\frac{nk}{2}}{\binom{n}{2}}$. Simplifying we obtain $p_{*}=\frac{k}{n-1}$. Thus, given $p_{*}$ and $n$, one needs to choose $k$ properly so that this criteria is satisfied.
3. Barabási-Albert model. The model, which was introduced by Albert and Barabási in 2002, is a random graph model [4] that uses preferential attachment (i.e. higher degree vertices tend to have even higher degrees) in order to generate scale-free networks. In NetworkX, samples from this model are generated via the barabasi_albert_graph(n, m, seed=None, initial_graph=None) function. The descriptions of the inputs in NetworkX references page are as follows.
  • $n$ int Number of nodes
  • $m$ int Number of edges to attach from a new node to existing nodes
In our case we use the default as the initial graph, and so begin with a star on $m+1$ nodes. Now, given the edge density $p_{*}$ and the number of nodes $n$, our goal is to select $m$ so that the resulting Barabási-Albert graph will have approximate edge density $p_{*}$.
Now, since we begin with $m+1$ nodes, the steps required to reach $n$ nodes is $N=n-(m+1)$. Then the number of edges inserted throughout these steps is $Nm=(n-(m+1))m$. Thus, also considering the initial nodes present, at the end of the evolution of the graph, we have
$(n-(m+1))m+m$
many edges, and so the edge density is
$\frac{(n-(m+1))m+m}{\binom{n}{2}}=\frac{nm-m^{2}}{\frac{n(n-1)}{2}}.$
Hence we would like to have $\frac{nm-m^{2}}{\frac{n(n-1)}{2}}\approx p_{*}$, or rearranging,
$2m^{2}-2nm+p_{*}nn(n-1)\approx 0$
Now the corresponding $\Delta$ is $4n^{2}-(4)(2)p_{*}n(n-1)=4n^{2}-8p_{*}n(n-1)$. Since $\Delta<0$ when $p_{*}$ is large for the number of nodes we intend to simulate, we will just consider the cases $p_{*}=0.1$ and $p_{*}=0.5$.
When $\Delta>0$, the root we use is given by
$m_{1}=\frac{2n-\sqrt{4n^{2}-8p_{*}n(n-1)}}{4}.$
For example, when $n=200$ and $p_{*}=0.1$, placing these in the given formula, we find $m_{1}=10.501$, which is to rounded to 11.
4. Random regular graphs. Random regular graphs are generated using random_regular_graph(d, n) function in NetworkX. Here, $n$ is the number of nodes and $d$ is the common degree. In this case, to achieve an edge density $p_{*}$, one needs to have $p_{*}=\frac{\frac{nd}{2}}{\binom{n}{2}}$, which can be rewritten as $p_{*}=\frac{d}{n-1}$. Thus, given $n$ and $p_{*}$, we choose $d$ appropriately in order to achieve the given edge density.
5.2 Results for the clustering index
Now we are ready to present the simulation results. In each fixed selection of parameters, 120-600 sample graphs of each model were generated for varying node counts, and these were averaged out to approximate the value of interest. For each plot, the number of nodes were chosen from the set $\{20,40,...380\}$, and the resulting averages were linearly interpolated. Note that the rewiring probability in Watts-Strogatz model does not affect the number of vertices nor the edge density; hence, in each of our simulations, we examined Watts-Strogatz graphs with rewiring probabilities 0.1, 0.3, 0.5, 0.7 and 0.9. In many of our simulations, Barabási-Albert graph and some of the Watts-Strogatz graphs had a few orders of magnitude larger clustering indices than the others; thus, we have created additional plots to better examine the remaining graph models.
Our expectation from our theoretical findings was that the growth of $\mathrm{CI}_{1}$ for Erdős-Rényi graphs was linear in $n$, which is verified in our simulation. One would also expect that, as the rewiring probability increases, the Watts-Strogatz graphs exhibit increasingly similar behavior to Erdős-Rényi graphs in terms of clustering, which seems to be the case.
We think that Barabási-Albert model having a large clustering index can be explained as follows: In the resulting graph after many iterations, the degrees of a significant portion of the nodes added to the initial star graph will be very small $(\leq 3)$. When the degree is small, the local clustering coefficient will be 0 or 1 with high probability. Hence, the clustering index can be expected to behave like $Cn^{2}$ asymptotically. This behavior seems to be more easily observed when the edge density increases.
Next we include the comparison of $\mathrm{CI}_{2}$ for the four models of interest. In the case of Erdős-Rényi, we showed that this index is upper bounded by a constant independent of the number of nodes, and conjectured that it is indeed also lower bounded by a constant. This is indeed the case according to our simulations, where $\mathrm{CI}_{2}$ can be observed to approach $\frac{2(1-p)(1-p^{2})}{p}$ as $n$ increases. Watts-Strogatz graphs with low rewiring probability and Barabási-Albert model again dominate the others in this case. We again observe that as the rewiring probability increases, $\mathrm{CI}_{2}$ of Watts-Strogatz model approaches that of Erdős-Rényi model.
It is also interesting to note that many of our random graphs seem to exhibit a peaking behavior around 40-60 nodes when edge density is 0.1. This behavior is most easily observed when edge density is low, and the position of peaks depend on the edge density; as edge density decreases, the number of nodes corresponding to the peak increases (up to a point). We think that the appearance of these peaks is a result of an equilibrium of two opposing effects:
When the edge density is 0.5, the Barabási-Albert model again dominates the others; this time by around 2 orders of magnitude at $n=380$. Watts-Strogatz model with lower rewiring probabilities again follow Barabási-Albert graphs in the ranking.
5.3 Results for the degree index
In this subsection we focus on the degree index and compare $\mathrm{DI}_{1}$ and $\mathrm{DI}_{2}$ for the same four random graph models for again varying values of the number of nodes. The first one below is for $\mathrm{DI}_{1}$ and the edge density in $\{0.1,0.5\}$.
For either of the edge densities considered, the Barabási-Albert model dominates the other ones significantly. This is not surprising since preferential attachment causes significant number of nodes to have degrees comparable to $n$ and a majority of the nodes to have very small degrees. This causes $\mathbb{E}[|d_{i}-d_{j}|]$ to be of order $n$, which then implies $\mathrm{DI}_{1}=\theta(n^{3})$.
We see that $\mathrm{DI}_{1}$ is zero for the case of random regular graphs, which is already trivial since the degrees in such graphs are the same. We can again observe that small world model behaves similarly to Erdős-Rényi as the rewiring probability increases.
Figure 6: Average $\mathrm{DI}_{2}$ values of the Erdős-Rényi graph, Barabási-Albert graph, random regular graph and Watts-Strogatz graphs with edge densities 0.1 and 0.5.
The above observations for $\mathrm{DI}_{1}$ are also true for $\mathrm{DI}_{2}$, with the modification that $\mathrm{DI}_{1}=\theta(n^{4})$ for Barabási-Albert, as expected. Note also that the $\mathrm{DI}_{2}$ values in our simulations are in agreement with our theoretical results.
6 Conclusion
In this manuscript, besides the well known degree index (or, irregularity), we introduced the concept of clustering index and studied the first moment of both of them in random graphs. Focusing on Erdős-Rényi graphs and the degree case, we were able to obtain either exact or asymptotic expressions for the expected degree indices of interest. Regarding the clustering index case, we obtained a linear upper bound for $\mathbb{E}[\mathrm{CI}_{1}(\mathcal{G})]$, and also showed that $\mathbb{E}[\mathrm{CI}_{2}(\mathcal{G})]$ is bounded above by a constant independent of $n$. Although we had theoretical and simulation based related analysis for random graphs, we have not done a detailed study of the indices of interest for real life complex networks, and we intend to do so in a subsequent work.
One main motivation for studying these indices in our case is their possible use in classification algorithms as features. Characteristics such as average clustering and the average degree are already used in various such problems, and we intend to do an experimental study on the use of degree and clustering indices to see whether they provide improvements in performance in classification problems. A second related path is in the field of finance, more particularly on the detection of financial crises. In a recent work [13], it is argued that a certain graph irregularity measure, namely the Laplace energy, can be used to detect economic crises. We plan to visit this issue with the indices discussed in current paper, and see whether the degree and clustering indices (as certain types of irregularities) can also be considered as a precursor of financial crises.
A last direction we are willing to follow is on computing the exact asymptotics of $\mathrm{CI}_{1}$ and $\mathrm{CI}_{2}$ in case of Erdős-Rényi graphs. For example, focusing on $\mathrm{CI}_{2}$, we were able to obtain a constant upper bound, and simulations suggested that this index indeed like a constant as the number of nodes increases. It would be interesting to verify the validity of this convergence theoretically.
@@ 6 Conclusion
Acknowledgements. Parts of this paper were completed at the Nesin Mathematics Village, the authors would like to thank Nesin Mathematics Village for their kind hospitality. The first author is supported by the BAGEP Award of the Science Academy, Turkey.

### v2

@@ 4 Degree index in Erdős-Rényi graphs
Remark 4.1
One may consider the normalized versions of the degree indices we study above in order to be able to do comparisons among distinct random graphs models in a clear way and to make meaningful inferences about the graph properties. Motivated by Theorem 4.1 and Theorem 4.2 above, a natural candidate for the scaled version of the statistic we study is
$\mathrm{DI}_{\alpha}^{*}=\frac{\mathrm{DI}_{\alpha}(\mathcal{G})}{n^{2+\alpha/2}}.$
Note that in the case of Erdős-Rényi graphs, as $n\rightarrow\infty$, $\mathbb{E}[\mathrm{DI}_{\alpha}^{*}]$ converges to $\sqrt{\frac{p(1-p)}{\pi}}$ and p(1p)p(1-p) for $\alpha$ equals 1 and 2, respectively.
Also, a second natural possible normalization for $\mathrm{DI}_{\alpha}(\mathcal{G})$ could be a scaling by $\binom{n}{2}$, where $n$ is the number of nodes. However, such a choice does not take $\alpha$ into account, and the scaled limits do not tend to constant values.
5 Monte Carlo simulations for other random graph models
The calculations for the degree and clustering indices become more involved when one leaves the framework of Erdős-Rényi random graphs and deals with other models. In this section, we focus on three other random graph models and do a Monte Carlo simulation study for these. The models we will be interested in are the Watts-Strogatz model, Barabási-Albert model and lastly, random regular graphs. The degree index in random regular graphs is already zero, and one is only interested in the clustering index for this special case.
5.1 Random graph models and implementation
Here, instead of providing a detailed technical background on the random graph models of interest, we will briefly discuss the practical aspects of the implementation of Monte Carlo simulations. In particular, we will focus on the selection of parameters so that the experiments are performed with matching edge densities for distinct random graph models. Before moving further, let us note that we used the NetworkX module in Python for all the simulations done below. Now, let us review the models under study, keeping in mind that we are willing to have a given edge density $p^{*}\in(0,1)$ in our simulations.
1. Erdős-Rényi model. The function erdos_renyi_graph which has parameters $n$ and $p$ as discussed in previous sections produces an Erdős-Rényi graph in NetworkX. One just chooses $p=p^{*}$ in order to satisfy the edge density criterion.
2. Watts-Strogatz model. The model was introduced by D. J. Watts and S. Strogatz in 1998, and it catches certain common properties of real-life complex networks by having the small world property, high clustering, and short average path lengths [25]. The description of watts_strogatz_graph function in NetworkX reference page is as follows: First create a ring over $n$ nodes. Then each node in the ring is joined to its $k$ nearest neighbors (or k1k-1 neighbors if is odd). Then shortcuts are created by replacing some edges as follows: for each edge $(u,v)$ in the underlying “$n$-ring with $k$ nearest neighbors” with probability $p$ (rewiring probability) replace it with a new edge $(u,w)$ with uniformly random choice of existing node $w$.
The number of edges in the Watts-Strogatz model is constant and is given by $\frac{nk}{2}$. Thus, in order to have the edge density $p^{*}$, one needs to have $p^{*}=\frac{\frac{nk}{2}}{\binom{n}{2}}$. Simplifying, we obtain $p^{*}=\frac{k}{n-1}$. Thus, given $p^{*}$ and $n$, one needs to choose $k$ properly so that this criterion is satisfied.
3. Barabási-Albert model. The model, which was introduced by Albert and Barabási in 2002, is a random graph model [4] that uses preferential attachment (i.e. higher-degree vertices tend to have even higher degrees) in order to generate scale-free networks. In NetworkX, samples from this model are generated via the barabasi_albert_graph(n, m, seed=None, initial_graph=None) function. The descriptions of the inputs in NetworkX references page are as follows.
  • $n$ int Number of nodes
  • $m$ int Number of edges to attach from a new node to existing nodes
In our case, we use the default as the initial graph and so begin with a star on $m+1$ nodes. Now, given the edge density $p^{*}$ and the number of nodes $n$, our goal is to select $m$ so that the resulting Barabási-Albert graph will have approximate edge density $p^{*}$.
Now, since we begin with $m+1$ nodes, the steps required to reach $n$ nodes is N=n(m+1)N=n-(m+1). Then the number of edges inserted throughout these steps is Nm=(n(m+1))mNm=(n-(m+1))m. Thus, also considering the initial nodes present, at the end of the evolution of the graph, we have
(n(m+1))m+m(n-(m+1))m+m
many edges, and so the edge density is
$\frac{(n-(m+1))m+m}{\binom{n}{2}}=\frac{nm-m^{2}}{\frac{n(n-1)}{2}}.$
Hence we would like to have $\frac{nm-m^{2}}{\frac{n(n-1)}{2}}\approx p^{*}$, or rearranging,
$2m^{2}-2nm+n(n-1)p^{*}\approx 0.$
Now the corresponding $\Delta$ is $4n^{2}-(4)(2)n(n-1)p^{*}=4n^{2}-8n(n-1)p^{*}$. Since $\Delta<0$ when $p^{*}$ is large for the number of nodes we intend to simulate, we will only consider the cases $p^{*}=0.1$ and $p^{*}=0.5$.
When $\Delta>0$, the root we use is given by
$m^{*}=\frac{2n-\sqrt{4n^{2}-8n(n-1)p^{*}}}{4}.$ (9)
For example, when $n=200$ and $p^{*}=0.1$, placing these in the given formula, we find $m^{*}=10.501$, which is rounded to 11. Note that for fixed $p^{*}$, $m^{*}$ depends linearly on $n$. In particular, when $p^{*}=0.5$, $m^{*}\approx\frac{n}{2}$ and when $p^{*}=0.1$, $m^{*}\approx\frac{n}{20}$.
4. Random regular graphs. Random regular graphs are generated using random_regular_graph(d, n) function in NetworkX. Here, $n$ is the number of nodes and $d$ is the common degree. In this case, to achieve an edge density $p^{*}$, one needs to have $p^{*}=\frac{\frac{nd}{2}}{\binom{n}{2}}$, which can be rewritten as $p^{*}=\frac{d}{n-1}$. Thus, given $n$ and $p^{*}$, we choose $d$ appropriately to achieve the given edge density.
5.2 Results for the clustering index
Now we are ready to present the simulation results. In each fixed selection of parameters, 120-600 sample graphs of each model were generated for varying node counts, and these were averaged to approximate the value of interest. For each plot, the number of nodes were chosen from the set $\{20,40,...,380\}$, and the resulting averages were linearly interpolated. Note that the rewiring probability in the Watts-Strogatz model does not affect the number of vertices nor the edge density; hence, in each of our simulations, we examined Watts-Strogatz graphs with rewiring probabilities of 0.1, 0.3, 0.5, 0.7 and 0.9. In many of our simulations, Barabási-Albert graph and some of the Watts-Strogatz graphs had growth rates significantly larger than the others; thus, we have created additional plots to better examine the remaining graph models in all our plots.
The following figure contains our results for $\mathrm{CI}_{1}$. Note that these plots are for the Barabási-Albert model where the parameter $m=m^{*}$ is chosen according to (9) in order to satisfy the edge density condition. In particular, this is in contrast to the standard Barabási-Albert model where $m$ is a fixed positive integer. One would naturally obtain a distinct growth rate for the standard case. We will have a more detailed relevant discussion below when we focus on the degree index.
Our expectation from our theoretical findings was that the growth of $\mathrm{CI}_{1}$ for Erdős-Rényi graphs was sublinear in $n$, which is verified in our simulation. One would also expect that, as the rewiring probability increases, the Watts-Strogatz graphs exhibit increasingly similar behavior to Erdős-Rényi graphs in terms of clustering, which seems to be the case.
For the Barabási-Albert model with $m=m^{*}$, when the edge density is 0.5, both $\mathrm{CI}_{1}$ and $\mathrm{CI}_{2}$ seem to be of order $n^{2}$. This is caused by the vastly different clustering behavior of the leaves of the initial star graph and the other nodes.  [^4] Since $m\approx n/2$ in this case, the initial star graph has $m-1\approx n/2-1$ leaves. Our simulation results show that, when n is large, the local clustering coefficients of these leaves are significantly higher compared to the clustering coefficients of the other nodes on average. This causes $\mathbb{E}[\mathrm{CI}_{1}]\geq K_{1}\binom{n}{2}$ and $\mathbb{E}[\mathrm{CI}_{2}]\geq K_{2}\binom{n}{2}$ for positive constants $K_{1},K_{2}$.
Next, in Figure 5, we include the comparison of $\mathrm{CI}_{2}$ for the four models of interest. In the case of Erdős-Rényi, we showed that this index is upper bounded by a constant independent of the number of nodes and we conjecture that it is indeed also lower bounded by a constant. This is indeed the case according to our simulations, where $\mathrm{CI}_{2}$ can be observed to approach $\frac{2(1-p)(1-p^{2})}{p}$ as $n$ increases. Watts-Strogatz graphs with low rewiring probability and the Barabási-Albert model again dominate the others in this case. Also, as in the case of $\mathrm{CI}_{1}$, we observe that as the rewiring probability increases, $\mathrm{CI}_{2}$ of the Watts-Strogatz model approaches that of Erdős-Rényi model.
It is also interesting in Figure 5 to note that many of our random graphs seem to exhibit a peaking behavior around 40-60 nodes when the edge density is 0.1. We think that the appearance of these peaks is a result of an equilibrium of two opposing effects:
When the edge density is 0.5, the Barabási-Albert model with $m=m^{*}$ again dominates the others; it seems to exhibit $\Theta(n^{2})$ growth while the others are sublinear. Watts-Strogatz model with lower rewiring probabilities again follows Barabási-Albert case in the ranking.
5.3 Results for the degree index
In this subsection, we focus on the degree index and compare $\mathrm{DI}_{1}$ and $\mathrm{DI}_{2}$ for the same four random graph models again for varying values of the number of nodes. Figure 6 is for $\mathrm{DI}_{1}$ and the edge density in $\{0.1,0.5\}$. Note that again there are two versions corresponding to each edge density due to the domination of Barabási-Albert model with $m=m^{*}$. In particular, the ones on the right do not contain the case for the Barabási-Albert model.
We see that as in the case of the clustering index, the small world model behaves similarly to Erdős-Rényi as the rewiring probability increases. In order to examine the behaviors for different models in more detail, we also include the $\log-\log$ plots for the graphs on the left in Figure 6.
Figure 7: $\log-\log$ plots for the $\mathrm{DI}_{1}$ values of the Erdős-Rényi graph, modified Barabási-Albert graph, and Watts-Strogatz graphs with edge densities 0.1 and 0.5.
The $\log-\log$ plot clearly indicates that the growth of $\mathbb{E}[\mathrm{DI}_{1}]$ for the preferential attachment model we consider is of higher order than the growth of other cases. The simulations we have and the heuristic arguments we use suggest that the growth rate in this case is of order $\Theta(n^{3})$, which is strictly larger than the one for Erdős-Rényi graphs (which is $\Theta(n^{5/2})$). However, note that the Barabási-Albert model we consider is different from the standard Barabási-Albert model in the following sense. Here, in order to have a fixed edge density for comparison among distinct models, the corresponding parameter $m$ increases as a function of $n$ as explained in Section 5.1; see (9). Thus, if we consider the standard Barabási-Albert model where $m$ is fixed, then we would obtain a different behavior. Indeed, we were informed by an anonymous reviewer that the maximal degree in the standard Barabási-Albert model is of order $\sqrt{n}$; see, for example, page 280 of [13]. From this, one immediately obtains a trivial upper bound of order $n^{5/2}$ for $\mathrm{DI}_{1}$ for fixed $m$. To have a better understanding, we also simulated the standard Barabási-Albert model. Figure 8 contains the corresponding $\log-\log$-plot.
Figure 8: $\log-\log$ plots for the $\mathrm{DI}_{1}$ and $\mathrm{DI}_{2}$ values of the Barabási-Albert graphs with $m\in\{1,2,3,5,10\}$, where $m$ denotes the number of edges to attach from a new node to existing nodes. The number of nodes is from 500 to 6000, with increments of 500.
The plots in Figure 8 and the related slope calculations suggest that $\mathbb{E}[\mathrm{DI}_{1}]=\Theta(n^{2})$ when $m$ is fixed. This seems natural since in this case the degrees of most nodes tend to be small and close to each other, which then causes $\mathrm{DI}_{1}$ to be small. On the other hand, in the Barabási-Albert model with $m=m^{*}$, (9) gives $m^{*}\sim Cn$ for some $C\in(0,1)$. In turn, this causes a significant number of nodes to have degrees comparable to $n$ and a remarkable number of nodes to have significantly smaller degrees, with the result $\mathbb{E}[|d_{i}-d_{j}|]$ being of order $n$, and $\mathbb{E}[\mathrm{DI}_{1}]$ of order $n^{3}$.
For the $\mathrm{DI}_{2}$ case, the simulation results are summarized in Figure 9. The observations here are similar to the $\mathrm{DI}_{1}$ case, and we do not detail these to avoid repetition.
Figure 9: Average $\mathrm{DI}_{2}$ values of the Erdős-Rényi graph, modified Barabási-Albert graph, random regular graph and Watts-Strogatz graphs with edge densities 0.1 and 0.5.
We close this section by noting that we also analyzed the $\log-\log$ plot for the case $\mathrm{DI}_{2}$ for the Barabási-Albert model with $m=m^{*}$, and that the related observations suggest $\mathbb{E}[\mathrm{DI}_{2}]=\Theta(n^{4})$.
6 Conclusion
In this manuscript, in addition to the well-known degree index (or, irregularity), we introduced the concept of clustering index and studied the first moment of both in a random graph framework. Focusing on Erdős-Rényi graphs and the degree case, we were able to obtain either exact or asymptotic expressions for the expected degree indices of interest. Regarding the clustering index case, we obtained a linear upper bound for $\mathbb{E}[\mathrm{CI}_{1}(\mathcal{G})]$, and also showed that $\mathbb{E}[\mathrm{CI}_{2}(\mathcal{G})]$ is bounded above by a constant independent of $n$. Although we had a theoretical and simulation-based related analysis for random graphs, we have not done a detailed study of the indices of interest for real-life complex networks and we intend to do so in a subsequent work.
One main motivation for studying these indices in our case is their possible use in artificial intelligence classification algorithms as features. Characteristics such as average clustering and the average degree are already used in various such problems, and we intend to do an experimental study in future work to see whether the use of degree and clustering indices provides improvements in performance in classification problems. A second related path is in the field of finance, more particularly in the detection of financial crises. In a recent work [14], it is argued that a certain graph irregularity measure, namely the Laplace energy, can be used to detect economic crises. Recently, we did some preliminary experiments to see whether the indices introduced in this paper could also be considered for such purposes. The initial analysis indicated that both are useful in this direction, but that the results for degree index are more statistically significant. We plan to conduct a rigorous analysis on whether degree and clustering indices (as certain types of network irregularities) can also be considered as a precursor of financial crises in upcoming work.
The last direction we are willing to follow is in computing the exact asymptotics of $\mathrm{CI}_{1}$ and $\mathrm{CI}_{2}$ in the case of Erdős-Rényi graphs. For example, focusing on $\mathrm{CI}_{2}$, we were able to obtain a constant upper bound, and simulations suggested that this index indeed behaves like a constant as the number of nodes increases. It would be interesting to verify this convergence theoretically.
Appendix: Proof of Lemma 4.1
@@ 6 Conclusion
(i) Let $X_{n},Y_{n}$ be independent binomial random variables with parameters $n$ and $1/2$. We claim that $\mathbb{E}|X_{n}-Y_{n}|=\frac{n\binom{2n}{n}}{2^{2n}}$. For this purpose, letting $Z_{n}$ be an independent binomial random variable with parameters $n$ and $1/2$, observe that
@@
(ii) Let $Z_{1},Z_{2}$ be independent standard normal random variables. We denote the Wasserstein distance between the distributions by $d_{W}$. Now, letting $\mu_{n}=\mathbb{E}[X_{n}]$ and $\sigma_{n}^{2}=\mathrm{Var}(X_{n})$, recall that
@@
Acknowledgments. We would like to thank the anonymous reviewers whose suggestions and corrections improved the paper significantly. Parts of this paper were completed at the Nesin Mathematics Village, the authors would like to thank Nesin Mathematics Village for their kind hospitality. The first author is supported by the BAGEP Award of the Science Academy, Turkey.